If a multiple choice test consist of 5 question each with four possible answers of which only one is correct , in how many different ways can a student check one of the question?
Found 2 solutions by stanbon, richwmiller:
Answer by stanbon(75887)

If a multiple choice test consists of 5 questions each with four possible answers of which only one is correct, in how many different ways can a student check one of the question?
----
# of ways to pick a question: 5
# of possible answers:: 4
-------------------------------
# of possible checks the student can make = 5*4 = 20
===============================================
Cheers,
Stan H.

Answer by richwmiller(17219)

I see at least three different ways to read this.
Is the question
1)"How many ways can a student answer each question?"
meaning say only question 6
4 ways per question
or
2)"How many ways can a student answer all of the questions?"
a)
1 way because he can only hand in one test with one set of answers
b) 5 because he can only hand in one test with one set of answers but he must choose 5 different answers
or
c)20 = 5*4 those are the choices before the student.
or
3) "How many different ways can the test be answered?"
4^5= 4*4*4*4*4=1024
There are 4 choices for 5 questions. And each choice is independent of the others.